yz(y+z)dx+zx(z+x)dy+xy(x+y)dz=0
1
Expert's answer
2020-09-21T16:14:17-0400

Given differential equation is

"yz(y+z)dx +xz(x+z)dy+xy(x+y)dz = 0"


Let "\vec{X} = yz(y+z)\hat{i} +xz(x+z)\hat{j}+xy(x+y)\hat{k}"

To check whether it is integral , "\vec{X} .(\nabla\times{\vec{X}}) = 0"

Now, "\nabla\times{\vec{X}} = \begin{vmatrix}
 \hat{i} & \hat{j} &\hat{k}\\
 \frac{\partial }{\partial x} & \frac{\partial }{\partial y} &\frac{\partial }{\partial z}\\
 yz(y+z) & xz(x+z) & xy(x+y)

\end{vmatrix}"



"\nabla\times{\vec{X}} = 2x(y-z )\hat{i} - 2y(x-z)\hat{j} + 2z(x-y)\hat{k}"

Now, "\vec{X} .(\nabla\times{\vec{X}}) =[ yz(y+z)\hat{i} +xz(x+z)\hat{j}+xy(x+y)\hat{k}
] .[2x(y-z )\hat{i} - 2y(x-z)\hat{j} + 2z(x-y)\hat{k}] = 0"

Hence it it intergrable.


Let "y = ux \implies dy = xdu + udx"

and "z = vx \implies dz = vdx+xdv"

Then partial differential equation will be reduced to the form

"uxvx(ux+vx)dx + vx^2(x+vx)(udx+xdu) + ux^2(x+ux)(vdx+xdv) = 0"


It can be further reduced to

"\frac{dx}{x} + \frac{du}{2u}+\frac{dv}{2v}-\frac{du}{2(u+v+1)}-\frac{dv}{2(u+v+1)} = 0"

Integrating it, we get


"\frac{x^2 uv}{u+v+1} = C"

putting values of u and v,


"xyz = C(x+y+z)"